If n is a positive integer, which of the following cannot be the units digit of 7^n ?
(A)1 (B) 3 (C) 5 (D) 7 (E) 9


Answer by JBarnum(2146)

5 cannot ever be in the units digit
all u have to do is use calculator
7^1=7
7^2=.9
7^3=..3
7^4=...1
7^5=....7
7^6=.....9
7^7=.....3
7^8=......1
7^9=.......7
7^10=........9
all the possitive numbers beyond will conclude to ending in only 1,3,7,9
you can see the pattern 7931 7931 79...